# Spectral decomposition and entropy for open set-valued maps

### v1

@@ 1. Introduction
One of the central problems in dynamical systems during the last century was the characterization of structurally stable systems. For many years, the so-called stability conjecture stimulated intense research activity; see, for instance, [peixoto1962structural] and [smale1967differentiable] for two landmark results. The conjecture was ultimately proved by Mañé [mane1987proof]. As a consequence of these and many subsequent contributions, it is now well understood that structural stability is deeply connected with hyperbolicity and, in particular, with the shadowing property of $\varepsilon$-orbits.
Recall that for a map $F:X\to X$ on a metric space, an $\varepsilon$-orbit is a sequence $(y_{n})_{n\geq 0}$ satisfying $\operatorname{dist}(F(y_{n}),y_{n+1})<\varepsilon$. This concept motivates the definition of the set-valued map $F_{\varepsilon}:X\to 2^{X}$, given by $F_{\varepsilon}(x)=B_{\varepsilon}(F(x))$, where $2^{X}$ denotes the collection of nonempty subsets of $X$, and $B_{\varepsilon}(x)\subset X$ is the open ball of radius $\varepsilon$ centered at $x$. In this way, $\varepsilon$-orbits of $F$ correspond precisely to true orbits of $F_{\varepsilon}$. Under mild assumptions, each $F_{\varepsilon}$ is an example of an open set-valued map, that is, a map $f:X\to 2^{X}$ whose graph $\operatorname{graph}(f)\subset X\times X$ is open and such that $f(x)$ is nonempty, open and connected for every $x\in X$. When $X$ is compact, every open set-valued map $f$ contains $F_{\varepsilon}$ for some map $F:X\to X$ and $\varepsilon>0$.
A natural question is whether the study of open set-valued maps yields meaningful information about the underlying dynamics. Given dynamical information on $F_{\varepsilon}$, what can be inferred about the behavior of $F$ itself? To illustrate that such maps capture relevant dynamical properties, let us recall a classical open problem in hyperbolic dynamics: the Anosov transitivity conjecture. It is not known whether every Anosov diffeomorphism is transitive (see [smale1967differentiable]), although this is widely believed to be the case. Indeed, the conjecture is true on infranilmanifolds [manning1974there], and to date, there are no known examples of Anosov diffeomorphisms defined on manifolds that are not infranilmanifolds. Recent progress toward this conjecture has been made in [BMelo15, Micena22, MicenaHertz22], as well as in the forthcoming works [BMelo25, Salgado25]. Interestingly, the analogous statement for flows is false [FrankWilliams80].
Regarding the transitivity of open set-valued maps, if $F$ is transitive, then $F_{\varepsilon}$ is transitive for all $\varepsilon>0$. However, the converse is generally false. In this paper, we prove that the converse does hold in the case of Anosov diffeomorphisms:
@@
Several authors have investigated the dynamical properties of set-valued maps. The definition of transitivity used in Theorem A follows that introduced in [luo2020shadowing]. Other related studies include [cordeiro2016continuum, duarte2006combinatorial, metzger2017topological, pilyugin2008shadowingcontractive, pilyugin2008shadowinghiperbolic]. In particular, [metzger2017topological] addresses topological stability for set-valued maps using techniques such as the shadowing property and inverse limit constructions, while [pilyugin2008shadowingcontractive, pilyugin2008shadowinghiperbolic] study shadowing from a different viewpoint. It is worth emphasizing that, among all these contributions, only [duarte2006combinatorial, duarte2015stability] specifically deal with open set-valued maps.
@@ 2. Preliminary
Let $x\in X$ and $y\in f(x)$. Since $\operatorname{graph}(f)$ is open, there exist open sets $A\ni x$ and $B\ni y$ in $X$ such that $A\times B\subset\operatorname{graph}(f)$. This implies that $B\subset f(a)$ for every $a\in A$. By induction, the same property holds whenever $y\in f^{n}(x)$ for some $n=1,2,\ldots$. Since $A$ and $B$ are open, there is $\delta$ such that $B_{\delta}(x)\subset A$ and $B_{\delta}(y)\subset B$. This concludes the proof. ∎
@@
Let $f:X\to 2^{X}$ be an open set-valued map. Then there exist a map $F:X\to X$ and a constant $\varepsilon_{0}>0$ such that
@@
For each $x\in X$, by Proposition 2.1 there exist a point $y_{x}\in f(x)$ and a number $\delta_{x}>0$ such that $B_{\delta_{x}}(y_{x})\subset f(y)$ for every $y\in B_{\delta_{x}}(x)$. The family $\{B_{\delta_{x}}(x)\}_{x\in X}$ forms an open cover of $X$. By compactness, there exists a finite subcover $\{B_{\delta_{x}}(x)\}_{x\in I}$. Define $\varepsilon_{0}=\min\{\,\delta_{x}:x\in I\,\}$, and construct a map $F:X\to X$ as follows: for each $z\in X$, choose an index $x\in I$ such that $z\in B_{\delta_{x}}(x)$ and set $F(z)=y_{x}$. Then, by construction,
@@
We claim that for every $x\in X$, either $x\in\tilde{A}$ or $x\in\tilde{B}$. Indeed, if there existed $x\in X$ such that $f(x)$ meets both $A$ and $B$, then $A\cup B\supset f(x)$ with $A\cap f(x)\neq\emptyset$ and $B\cap f(x)\neq\emptyset$, contradicting the connectedness of $f(x)$. Hence $X=\tilde{A}\cup\tilde{B}$.
@@
$\Omega(f)=\{x\in X:x\rightsquigarrow x\}\qquad\text{and}\qquad\Omega_{final}(f)=\{x\in X:\mbox{ if }x\rightsquigarrow y,\mbox{ then }y\rightsquigarrow x\}.$
@@
For any open set-valued map $f$, one always has $\Omega_{\mathrm{final}}(f)\subset\Omega(f)$. Figure 1 illustrates the graph of an open set-valued map on $X=[0,1]$ for which $\Omega_{\mathrm{final}}(f)\neq\Omega(f)$. In this example, the points $x$ and $y$ belong to $\Omega(f)$; however, $y\notin\Omega_{\mathrm{final}}(f)$ since $y\rightsquigarrow x$ but $x\not\rightsquigarrow y$.
@@
We refer to $x]_{\Omega_{\mathrm{final}}}$ as a final class. There are only finitely many such classes, and each of them is an invariant set (see Theorem 3.3).
@@
Let $f$ be an open set-valued map. If $x\in\Omega_{final}$ then $x]_{\Omega}=[x]_{\Omega_{final}}$. In particular, both $x]_{\Omega_{final}}$ and $\Omega_{final}(f)$ are open sets.
@@
Let $x\in\Omega_{final}$. By definition, we have $x]_{\Omega_{final}}\subset[x]_{\Omega}$. On the other hand, given $y\in[x]_{\Omega}$, then $x\leftrightsquigarrow y$. If $y\rightsquigarrow z$ then $x\rightsquigarrow z$ by transitivity. Since $x\in\Omega_{final}$, we also have $z\rightsquigarrow x$. By transitivity again, $z\rightsquigarrow y$ which implies that $y\in\Omega_{final}$. Hence, $x]_{\Omega_{final}}=[x]_{\Omega}$. ∎
@@
For any open set-valued map $f:X\to 2^{X}$, the sets $\Omega(f)$ and $\Omega_{\mathrm{final}}(f)$ are always nonempty. Indeed, to see that $\Omega(f)\neq\emptyset$, consider an orbit $(x_{n})_{n\geq 0}$ of $f$. By compactness of $X$, there exists a subsequence such that $x_{n}\to x$ for some $x\in X$. By Propositions 2.1 and 2.2, for sufficiently large $n$ we have $x_{n}\rightsquigarrow x$ and $x\rightsquigarrow x_{n}$, so $x\in\Omega(f)$.
@@
Let $U,V\subset[x]_{\Omega_{\mathrm{final}}(f)}$ be nonempty open sets. Take $u\in U$ and $v\in V$. Since $u\rightsquigarrow v$, there exists $n\in\mathbb{N}$ such that $v\in f^{n}(u)$. Hence $f^{n}(U)\cap V\neq\emptyset$, which proves that the restriction $\left.\kern-1.2ptf\vphantom{\big|}\right|_{[x]_{\Omega_{\mathrm{final}}(f)}}$} is transitive. ∎
@@ 3. Spectral decomposition for open set-valued maps
$f(x)=\begin{cases}[0,1],&\text{if }x>2,\\[3.0pt] [3,4],&\text{if }x<2.\end{cases}$
@@
Let $X$ be a connected space, and let $f:X\to 2^{X}$ be a transitive open set-valued map. Suppose $U_{0}\subset X$ is an open set such that $\overline{U_{0}}\subset f(X)$ and $U_{0}\subset f^{n}(U_{0})$ for some $n\geq 1$. For each $k\geq 1$, define
@@
By Proposition 2.3, $f(X)$ is connected, so at least two of these sets intersect. Hence there exist distinct $a,b\in\{1,\dots,n\}$ such that
@@
Among the indices $\{k+m,k+m+1,\dots,k+m+n-1\}$ there is at least one multiple of $n$; denote it by $jn$. Therefore $x\in U_{jn}\subset V_{0}$, and the result follows. ∎
@@
Moreover, because $U\subset f^{n}(U)$, it follows that $f^{kn}(U_{j})\cap V\neq\emptyset$ for all $k\geq n_{j}$.
Let $N=\max\{n_{j}:0\leq j\leq n-1\}$. If m>Nn+(n1)m>Nn+(n-1), then there exists $j\in\{0,1,\dots,n-1\}$ and $k\in\mathbb{N}$ such that mj=knm-j=kn with $k\geq n_{j}$. Hence,
@@ 4. Topological entropy for open set-valued maps
In this section, we study the topological entropy of open multivalued maps. The notion we adopt is parallel to the classical definition of topological entropy for univalued maps and was developed in [cordeiro2016continuum]. For completeness, we briefly recall the basic definitions and refer to these references for additional details.
@@
$(x_{0},\dots,x_{n-1}),\ (y_{0},\dots,y_{n-1})\in E,$
there exists $0\leq i\leq n-1$ such that $\operatorname{dist}(x_{i},y_{i})\geq\varepsilon$. We denote by $s_{n}(\varepsilon)$ the maximal cardinality of an $(n,\varepsilon)$-separated subset of $\operatorname{Orb}_{n}(f)$.

### v2

@@ 1. Introduction
One of the central problems in dynamical systems during the last century was the characterization of structurally stable systems. For many years, the so-called stability conjecture stimulated intense research activity; see, for instance, [peixoto1962structural] and [smale1967differentiable] for two landmark results. The conjecture was ultimately proved by Mañé [mane1987proof]. As a consequence of these and many other contributions, it is now well understood that structural stability is deeply connected to hyperbolicity and, in particular, to the shadowing property of $\varepsilon$-orbits.
A discrete dynamical system is defined by a map $F:X\to X$ on a metric space. The orbits of $F$, which describe the evolution of points in $X$, are the sequences of iterates $(x_{n})_{n\geq 0}$, where $x_{n}=F^{n}(x_{0})$. An $\varepsilon$-orbit is a sequence $(y_{n})_{n\geq 0}$ satisfying $\operatorname{dist}(F(y_{n}),y_{n+1})<\varepsilon$. We say that the orbit $(x_{n})$ $\delta$-shadows the $\varepsilon$-orbit $(y_{n})$ if $\operatorname{dist}(x_{n},y_{n})<\delta$. These notions motivate the definition of the set-valued map $F_{\varepsilon}:X\to 2^{X}$, given by $F_{\varepsilon}(x)=B_{\varepsilon}(F(x))$, where $2^{X}$ denotes the collection of nonempty subsets of $X$, and $B_{\varepsilon}(x)\subset X$ is the open ball of radius $\varepsilon$ centered at $x$. Then $\varepsilon$-orbits of $F$ correspond to true orbits of $F_{\varepsilon}$. Under mild assumptions, each $F_{\varepsilon}$ is an example of an open set-valued map, that is, a map $f:X\to 2^{X}$ whose graph $\operatorname{graph}(f)\subset X\times X$ is open and such that $f(x)$ is nonempty, open, and connected for every $x\in X$. When $X$ is compact, every open set-valued map $f$ contains $F_{\varepsilon}$ for some map $F:X\to X$ and $\varepsilon>0$.
A natural question is whether the study of open set-valued maps provides meaningful information about the underlying dynamics. Given dynamical information on $F_{\varepsilon}$, what can be inferred about $F$’s own behavior? To illustrate that such maps capture relevant dynamical properties, let us recall a classical open problem in hyperbolic dynamics: the Anosov transitivity conjecture. It remains unknown whether every Anosov diffeomorphism is transitive (see [smale1967differentiable]), although this is widely believed to be the case. The conjecture is true on infranilmanifolds [manning1974there], and so far, no examples of Anosov diffeomorphisms defined on manifolds that are not infranilmanifolds have been found. Recent advances toward this conjecture appear in [BMelo15, Micena22, MicenaHertz22], and in the forthcoming works [BMelo25, Salgado25]. Interestingly, the analogous statement for flows is false [FrankWilliams80].
Regarding the transitivity of open set-valued maps, if $F$ is transitive, then $F_{\varepsilon}$ is transitive for all $\varepsilon>0$. However, the converse is generally false. In this paper, we prove that the converse holds in the case of Anosov diffeomorphisms:
@@
Several authors have investigated the dynamical properties of set-valued maps. The definition of transitivity used in Theorem A follows that introduced in [luo2020shadowing]. Other related studies include [cordeiro2016continuum, duarte2006combinatorial, metzger2017topological, pilyugin2008shadowingcontractive, pilyugin2008shadowinghiperbolic]. In particular, [metzger2017topological] addresses topological stability for set-valued maps using techniques such as the shadowing property and inverse limit constructions, while [pilyugin2008shadowingcontractive, pilyugin2008shadowinghiperbolic] studies shadowing from a different viewpoint. It is worth emphasizing that, among all these contributions, only [duarte2006combinatorial, duarte2015stability] specifically deals with open set-valued maps.
@@ 2. Preliminary
Let $x\in X$ and $y\in f(x)$. Since $\operatorname{graph}(f)$ is open, there exist open sets $A\ni x$ and $B\ni y$ in $X$ such that $A\times B\subset\operatorname{graph}(f)$. This implies that $B\subset f(a)$ for every $a\in A$. By induction, the same property holds when $x\in X$ and $y\in f^{n}(x)$. Since $A$ and $B$ are open, there exists $\delta>0$ such that $B_{\delta}(x)\subset A$ and $B_{\delta}(y)\subset B$. This concludes the proof. ∎
@@
Let $f:X\to 2^{X}$ be an open set-valued map. Then there exists a map $F:X\to X$ and a constant $\varepsilon_{0}>0$ such that
@@
For each $x\in X$, by Proposition 2.1, there exists a point $y_{x}\in f(x)$ and a number $\delta_{x}>0$ such that $B_{\delta_{x}}(y_{x})\subset f(y)$ for every $y\in B_{\delta_{x}}(x)$. The family $\{B_{\delta_{x}}(x)\}_{x\in X}$ forms an open cover of $X$. By compactness, there exists a finite subcover $\{B_{\delta_{x}}(x)\}_{x\in I}$. Define $\varepsilon_{0}=\min\{\,\delta_{x}:x\in I\,\}$, and construct a map $F:X\to X$ as follows: for each $z\in X$, choose an index $x\in I$ such that $z\in B_{\delta_{x}}(x)$ and set $F(z)=y_{x}$. Then, by construction,
@@
We claim that for every $x\in X$, either $x\in\tilde{A}$ or $x\in\tilde{B}$. Indeed, if there exists $x\in X$ such that $f(x)$ meets both $A$ and $B$, then $A\cup B\supset f(x)$ with $A\cap f(x)\neq\emptyset$ and $B\cap f(x)\neq\emptyset$, contradicting the connectedness of $f(x)$. Hence $X=\tilde{A}\cup\tilde{B}$.
@@
$\Omega(f)=\{x\in X:x\rightsquigarrow x\}\qquad\text{and}\qquad\Omega_{\mathrm{final}}(f)=\{x\in X:\mbox{ if }x\rightsquigarrow y,\mbox{ then }y\rightsquigarrow x\}.$
@@
For any open set-valued map $f$, one always has $\Omega_{\mathrm{final}}(f)\subset\Omega(f)$. Figure 1 illustrates the graph of an open set-valued map on $X=[0,1]$ for which $\Omega_{\mathrm{final}}(f)\neq\Omega(f)$. In this example, the points $x$ and $y$ belong to $\Omega(f)$; however, $y\notin\Omega_{\mathrm{final}}(f)$ since $y\rightsquigarrow x$, but $x\not\rightsquigarrow y$.
@@
We refer to each $x]_{\Omega_{\mathrm{final}}}$ as a final class. There are only finitely many such classes, and each of them is an invariant set (see Theorem 3.3).
@@
Let $f$ be an open set-valued map. If $x\in\Omega_{\mathrm{final}}$, then $x]_{\Omega}=[x]_{\Omega_{\mathrm{final}}}$. In particular, both $x]_{\Omega_{\mathrm{final}}}$ and $\Omega_{\mathrm{final}}(f)$ are open sets.
@@
Let $x\in\Omega_{\mathrm{final}}$. By definition, we have $x]_{\Omega_{\mathrm{final}}}\subset[x]_{\Omega}$. On the other hand, given $y\in[x]_{\Omega}$, then $x\leftrightsquigarrow y$. If $y\rightsquigarrow z$, then $x\rightsquigarrow z$ by transitivity. Since $x\in\Omega_{\mathrm{final}}$, we also have $z\rightsquigarrow x$. By transitivity again, $z\rightsquigarrow y$, which implies that $y\in\Omega_{\mathrm{final}}$. Hence, $x]_{\Omega_{\mathrm{final}}}=[x]_{\Omega}$. ∎
@@
For any open set-valued map $f:X\to 2^{X}$, the sets $\Omega(f)$ and $\Omega_{\mathrm{final}}(f)$ are always nonempty. Indeed, to see that $\Omega(f)\neq\emptyset$, consider an orbit $(x_{n})_{n\geq 0}$ of $f$. By the compactness of $X$, there exists a subsequence such that $x_{n}\to x$ for some $x\in X$. By Propositions 2.1 and 2.2, for sufficiently large $n$, we have $x_{n}\rightsquigarrow x$ and $x\rightsquigarrow x_{n}$, so $x\in\Omega(f)$.
@@
Let $U,V\subset[x]_{\Omega_{\mathrm{final}}(f)}$ be nonempty sets. Take $u\in U$ and $v\in V$. Since $u\rightsquigarrow v$, there exists $n\in\mathbb{N}$ such that $v\in f^{n}(u)$. Hence $f^{n}(U)\cap V\neq\emptyset$, which proves that the restriction of $f$ to $x]_{\Omega_{\mathrm{final}}(f)}$ is transitive. ∎
@@ 3. Spectral decomposition for open set-valued maps
$f(x)=\begin{cases}[0,1],&\text{if }x\in[3,4]\\[3.0pt] [3,4],&\text{if }x\in[0,1].\end{cases}$
@@
Let $X$ be a connected space, and let $f:X\to 2^{X}$ be a transitive open set-valued map. Suppose $U_{0}\subset X$ is an open set such that $U_{0}\subset f^{n}(U_{0})$ for some $n\geq 1$. For each $k\geq 1$, define
@@
By Proposition 2.3, $f(X)$ is connected, so at least two of these sets intersect. Hence, there exist distinct $a,b\in\{1,\dots,n\}$ such that
@@
Among the indices $\{k+m,k+m+1,\dots,k+m+n-1\}$, there is one multiple of $n$; denote it by $jn$. Therefore $x\in U_{jn}\subset V_{0}$, and the result follows. ∎
@@
Moreover, because $U_{0}\subset f^{n}(U_{0})$, it follows that $f^{kn}(U_{j})\cap V\neq\emptyset$ for all $k\geq n_{j}$.
Let $N=\max\{n_{j}:0\leq j\leq n-1\}$. If m>Nn+(n1)m>Nn+(n-1), then there exist $j\in\{0,1,\dots,n-1\}$ and $k\in\mathbb{N}$ such that mj=knm-j=kn with $k\geq n_{j}$. Hence,
@@ 4. Topological entropy for open set-valued maps
In this section, we study the topological entropy of open set-valued maps. The notion we adopt is parallel to the classical definition of topological entropy for single-valued maps and was developed in [cordeiro2016continuum]. For completeness, we briefly recall the basic definitions and refer to these references for additional details.
@@
$(x_{1},\dots,x_{n}),\ (y_{1},\dots,y_{n})\in E,$
there exists $1\leq i\leq n$ such that $\operatorname{dist}(x_{i},y_{i})\geq\varepsilon$. We denote by $s_{n}(\varepsilon)$ the maximal cardinality of an $(n,\varepsilon)$-separated subset of $\operatorname{Orb}_{n}(f)$.